OpenStudy (anonymous):

Aimee and b solved the system below 2x-3y=8 -2x+3y=-8 Aimee says the solution has no answer but b says their are infinite solutions. Who is correct? Justify

OpenStudy (texaschic101):

2x - 3y = 8 -2x + 3y = -8 -------------add 0 + 0 = 0 0 = 0 (when it comes out equal, this means their are infinite solutions. If it were to come out like 0 = 4, or 5 = 8, then there would be no solutions. So b is correct in saying infinite solutions.